# Economic DAO Governance: A Contestable Control Approach

### v3

@@ 4.1.3 Bidding Intensity, Group Bidding, and Collusion
Bidding intensity and the viability of the mechanism itself depends on capital market adequacy, the depth and development of capital markets. Making a bid through the mechanism requires deposits of the same order of magnitude as the total pre-bid value of the DAO. For established DAOs such as Uniswap, Maker, or Compound, deposit amounts of a billion dollars or more may be required. [^41] As of May 22, 2024, these three DAO governance tokens had market capitalizations of approximately \$5.5 billion, \$2.7 billion, and \$500 million respectively. Coinmarketcap (2024). A deep bench of institutional investors such as venture capital firms, private equity firms, and hedge funds is required to make the auction mechanism robust and create bidding intensity. At least on the venture capital front, a large number of such firms already exist and are very active with respect to cryptocurrency projects including many DAOs. [^42] Crowdcreate (2021). These firms would be a ready source of funding for auction bidders. The temporary contestable control created by the auction mechanism, if implemented, might itself spawn specialist institutions similar to hedge funds in the current corporate landscape that combine portfolio investment with selective activism in governance. [^43] Yin and Zhu (2023) provide an empirical analysis. Such hedge funds typically engage in activism with respect to only a small portion of their portfolios. Burkart and Lee (2021) summarize the evidence that takeover activism aimed at or resulting in changes in corporate control generate higher target and activist returns.
@@ 4.1.4 Toehold Reporting and Post-auction Market Dangers
Second, the true value of the toehold proportion, $t_{b}$, is necessary to enforce the market liquidity condition $t_{b}+t_{f}\leq t_{m}$, which guarantees that at least $t_{m}q$ tokens are held by others, creating a pool of market liquidity. Relevant to both purposes, nothing in the bidding mechanism itself creates an incentive for a true report, and, in fact, the incentive is to cheat by hiding some of the toehold to artificially inflate the bid, $A$.
@@
Could a bounty system be effective and code feasible? DAO positions appear as a set of public addresses and token quantities. Commercially available technologies to trace and attribute token ownership exist and have a significant degree of effectiveness at what may be feasible cost. [^49] See Yaffe-Bellany (2023) (discussing Chainalysis). But reliance on outside commercial parties inhibits code feasibility. The DAO may have to implement the bounty system via contracts with outside teams. Doing so effectively while the control party has control of the DAO may not be possible. The task is complicated by the possible need to resolve disputes about the veracity of identifications claimed by the bounty hunters. [^50] Dispute resolution is more difficult when the parties do not have accessible on-chain assets at risk. See note 44 supra. An approach such as requiring good faith deposits from bounty hunters might be required to implement a dispute resolution mechanism that is code feasible. And possible hidden ownership presents a major challenge to the effectiveness of bounty systems. [^51] Bounty hunters do have some possible strategies in the face of hidden ownership. Counterparties to the hidden ownership position used for concealment have a big incentive to collect the bounty by disclosing the failure to report, presumably earning an additional token at the expense of the control party without violating the underlying contract. Bounty hunters may angle for a cut by advertising this opportunity broadly and offering assistance. Although a bounty system may be too difficult to implement, it is potentially a very powerful deterrent. Successful bounty recovery is a disaster for the winning bidder, involving loss not only of potential added token value, the motivation for concealment, but also of the base value, $P_{0}$.
One set of approaches to address control party position reporting issues centers around using know your customer registration systems that make token holders identifiable. Requiring registration of all token positions at all times would be very costly both initially and in the face of continual revisions as tokens are traded between parties. Registration might also raise privacy considerations requiring costly zero-knowledge proof or other technologies to make identities private yet verifiable. Nonetheless, use of registration limited to subsets of token holders and to particular points in time might be useful and cost-effective as we describe in what follows.
@@
    Registration. Parties can register through a know your customer process to demonstrate that they are not the control party or related to the control party. [^52] Registration can be made consistent with privacy through approaches such as requiring a zero-knowledge proof confirmation that they are not a restricted party linked to the winning bidder. For example, Rosenberg et al. (2022) describe a zero-knowledge proof credentialing system that they call “zk-creds.” In one instantiation, parties embed their passport in a zero-knowledge privacy layer that allows proof of identity. All registered parties are eligible to participate in the token auction. The set of $T_{1}$ token holders who register will be eligible to receive surplus from the auction.
@@ 4.1.6 Auction Efficiency and Market Liquidity
Like second price auction frameworks in other contexts, the English auction that embodies the basic mechanism here has strong efficiency properties. The bidder who can produce the most social surplus wins, with appropriation of that surplus by the winning bidder limited by the level of the second highest bid. As discussed, there is another limitation: The mechanism is subject to the market liquidity constraint. This constraint limits the added token value available to the winning bidder, reducing it by $F=(1-t_{m})(V^{*}-P_{0})$, the minimum amount that must accrue to free riders. Consider the project $(V^{*},C^{*})$ that creates the highest possible social surplus $\psi^{*}=N^{*}-C^{*}$ and produces total added token value $N^{*}=(V^{*}-P_{0})q$. If $\psi^{*}=N^{*}-C^{*}>0>N^{*}-C^{*}-F$, then this best possible project will not be implemented because it cannot cover its cost and also distribute $F$ to the free riders.
@@
There is another possible source of inefficiency arising from the auction mechanism here, which we will call “toehold overbidding.” Burkart (1995) shows that in an English auction, it may be optimal for participants with toeholds to bid higher than their valuations. Overbids create a danger of overpaying if the overbidder wins, but they also may push up the winning auction price, benefiting the toehold position if the overbidder loses. Burkart (1995) proves that, in the context of his model, for sufficiently small overbids, the potential toehold benefit outweighs the danger of loss from winning the auction with an overbid. [^54] The potential gain on the toehold is of the order of magnitude of the overbid while the potential loss from the overbid is of the order of magnitude of the overbid squared. Burkart (1995) assumes a private values setting in his main inquiry. Bulow et al. (1999) show that overbidding tends to be a much more serious problem in common values situations. Note also that the potential loss of efficiency from overbidding is a function of bidding intensity. If the bidders are dense in the possible valuations up the maximum one, then the winning bidder will be pushed up to its maximum value and will not receive any of the social surplus generated by the project. Overbidding does not have any benefit. Burkart (1995) developed his results in a setting with only two bidders, a very low bidding intensity with a subsequent high likelihood of a substantial gap in valuations between the two bidders. Under the mechanism here, successful overbidding would induce the winning bidder to reduce its surplus claim in order to prevail in the auction, the reduction accruing in part to the overbidder. Successful overbidding is socially desirable because more social surplus is shifted to the $T_{1}$ token holders, which results in a positive investment impact while not affecting the result that the best business plan is put into effect. [^55] Recall that a smaller surplus claim reduces the proportion of $T_{1}$ token holdings that must be sold to the winning bidder at $P_{0}$ and increases the corresponding proportion on which the $T_{1}$ token holders can earn the full amount of added token value. Note that the strategies of raising the value claim and lowering the surplus claim are equivalent for the overbidder. An overbidder facing a break-even bid level of $A^{*}$ expects to lose the full amount of the overbid $A-A^{*}>0$ in either case. Unsuccessful overbidding, however, has negative social properties. The winning bidder implements an inferior business plan. To the extent of the overbid, additional apparent social surplus flows to the $T_{1}$ token holders based on the value deposit forfeit function, which will result in a positive investment impact. However, this impact is larger than is optimal, because at least part of the surplus is apparent rather than real, being in excess of what would be generated even by the best business plan. [^56] Investment is over-encouraged if we look at this auction event in isolation. It is possible that the transfer in excess of actual social surplus generated by this type of event offsets the failure to shift the full amount of social surplus to the $T_{1}$ token holders in other auctions. In short, it is possible that the transfer in excess of social surplus is second-best optimal in some instances.
The mechanism here makes toehold overbidding less profitable. In the absence of the mechanism, consider a bidder with a toehold proportion $t_{b}$ and project valuation $V$. In the case of two bidders considered in Burkart (1995), a bid of $X>V$ will result in a gain on the toehold of $t_{b}(X-V)$ if the other bid is higher than $X$ and a loss of the same magnitude if the other bid is less than $X$. Under the mechanism, however, the loss will include forfeit of part of the value deposit, with the result that the loss will be at least $(X-V)q$ while the potential toehold gain remains at $t_{b}(X-V)q$. Whether this difference will entirely preclude toehold overbidding depends on assumptions about the distribution of possible project values for each bidder and about whether auction participants are aware of the distributions of other participants. But it is clear that the mechanism discourages toehold overbidding compared to the status quo. [^57] If the value deposit is increased above the lower bound of $(1-t_{d})(S-P_{0})q$ with, say, a uniform percentage increase in the deposit forfeit amount, $\phi(D_{v},P_{0},S,X)$, for all outcomes $X\in[P_{0},S]$, then the potential loss from toehold overbidding will be even higher, discouraging toehold overbidding even more. However, increases in the value deposit and the deposit forfeit amounts above their lower bounds has other effects. In particular effort by control parties may be over-incentivized. See subsection 5.4 infra.
@@ 5.1 Impact on Investment
Another consideration is that sealed-bid first price auctions may avoid the collusion problems that can arise in second price auctions. Collusion coalition parties in a sealed-bid first price auction can defect by bidding above an agreed upon low price chosen to enable the highest valuing bidder to prevail at that price and thus collect the largest possible amount of added token value for the coalition. In a second price auction, the highest valuing bidder will submit their best bid, and other coalition members cannot win the auction by deviating. Deviating only reduces the joint added token value for the coalition. The potential considerations in choosing among auction forms are myriad, and we do not attempt a detailed discussion here. [^65] A good starting point is Krishna (2010). Instead, we simply rest on a claim that given current knowledge, the English auction is a strong candidate with respect to efficiency and raising revenue compared to other standard auctions.666666 We have been assuming a “private values” setting with asymmetric information in which bidder business plans are known only to each bidder and bidder valuations of the plans are independent. See note 40 supra. If, instead, there is a common values element in which bidders have different signals concerning the value of the current operation and bid based on perceived undervaluation at price $P_{0}$, then the efficiency and surplus-distribution of various auction methods are potentially different than under the private values assumption. However, the English auction performs relatively well in a common values setting. Under certain assumptions such as the average crossing property, the English auction remains efficient in the sense that there is an ex-post equilibrium in which the bidder who can add the most value will win. (Krishna, 2010, pp. 134-143). In this setting, the English auction tends to raise more revenue than other approaches, including sealed-bid first price and second price auctions. (Krishna, 2010, pp. 97-100). In the context here, raising more revenue equates to extracting more added token value from the winning bidder, one of our desiderata.
@@ 5.3 Undervaluation and Treasury Raid Scenarios
Potential “treasury raid” scenarios provide an interesting perspective on how the mechanism would operate. DAOs typically have treasuries consisting of cryptocurrency tokens that can be used for further development of the DAO. Treasuries can store retained earnings as well as funds received from investors. Treasuries may be under the direct control of token holders who can direct their use by voting on proposals. [^72] In some cases, there are other parties who have that formal power or a veto on token holder proposals for spending treasury resources. In one common organizational structure, the DAO is embodied as a Swiss Foundation. The Foundation acts a wrapper, creating legal personhood for the DAO, with concomitant limited liability. Under Swiss law the Foundation must have formal legal control over the DAO treasury. Although the Foundation may in practice typically defer to token holder votes, there is the power to veto vote outcomes if they involve disbursing treasury funds. Direct control by token holders creates the danger of “treasury raids.” The most pernicious example of a treasury raid is one in which a party gains control of the DAO and then transfers the entire treasury to itself by voting positively on a suitable proposal. The party is stealing the other token holders’ stakes, and the term “raid” along with its negative connotations is apt.
@@
First, there may be an expectation that the parties currently guiding the DAO, although good-intentioned, will engage in foolish business plans, gradually wasting the treasury assets. The foolish business plans may be the result of the exhaustion of the original guiding purposes of the DAO. In this case, the social optimum is to end the DAO and distribute the treasury to the token holders, permitting re-investment of the treasury assets in more promising projects. In a situation of capital market adequacy, there will be “vulture investors,” who facilitate this result via a takeover or through activism in the face of reluctant guiding or control parties.737373 Vulture investors already exist in the cryptocurrency space. The saga of Rook DAO, recounted by Gilbert, Aleksandar (2023) and Nelson (2023), is an example. Rook DAO encountered an internal conflict and was trading below the value of its treasury. Co-founders and members of the core team proposed shifting 75% of the tokens to themselves to continue the project, leaving 25% to the rest of the token community. Activist investors facilitated an increase in the community’s share to 60%. Liu (2023) details the claimed role of Arca, one of the activist investors. As described in Arca (2023) and Dorman (2020), Arca’s operations include some that resemble a hedge fund that both makes portfolio investments and engages in activism.
@@
If, instead, the token value was underwater because the guiding parties have a superior business plan that requires secrecy, capital market adequacy means that these parties can obtain funding in confidence to outbid parties intent on liquidation. The other token holders receive at least their pro rata share of the treasury and possibly also some of the social surplus from the envisioned superior business plan. If the guiding parties think they have a superior business plan but cannot convince any funding party of that opinion, then upon liquidation, the guiding parties at least receive their share of the treasury, possibly amounting to enough to implement their business plan through a new project. [^75] The outcome in the Rook DAO events described in note 73 supra appears to represent these kinds of splits between parties intent on implementing a business plan and parties who want to exit through a liquidation that brings the DAO value up to the level of the treasury. The Rook DAO treasury assets were divided between two such groups of parties. The bidding competition central to the mechanism creates a way to divide the treasury assets more cogently, with the key question being whether the parties who want to continue can outbid the liquidation value. If they can, the other parties receive at least the pro rata treasury value of their tokens and likely more. If they cannot, they receive their share of the treasury and can pursue the project with those funds.

### v4

@@ 4.1.3 Bidding Intensity, Group Bidding, and Collusion
Bidding intensity and the viability of the mechanism itself depends on capital market adequacy, the depth and development of capital markets. Making a bid through the mechanism requires deposits of the same order of magnitude as the total pre-bid value of the DAO. For established DAOs such as Uniswap, Maker, or Compound, deposit amounts of a billion dollars or more may be required. [^41] As of November 24, 2024, these three DAO governance tokens had market capitalizations of approximately \$6.4 billion, \$1.5 billion, and \$580 million respectively. Coinmarketcap (2024). A deep bench of institutional investors such as venture capital firms, private equity firms, and hedge funds is required to make the auction mechanism robust and create bidding intensity. At least on the venture capital front, a large number of such firms already exist and are very active with respect to cryptocurrency projects including many DAOs. [^42] Crowdcreate (2021). These firms would be a ready source of funding for auction bidders. The temporary contestable control created by the auction mechanism, if implemented, might itself spawn specialist institutions similar to hedge funds in the current corporate landscape that combine portfolio investment with selective activism in governance. [^43] Yin and Zhu (2023) provide an empirical analysis. Such hedge funds typically engage in activism with respect to only a small portion of their portfolios. Burkart and Lee (2021) summarize the evidence that takeover activism aimed at or resulting in changes in corporate control generate higher target and activist returns.
@@ 4.1.4 Toehold Reporting and Post-auction Market Dangers
Second, the true value of the toehold proportion, $t_{b}$, is necessary to enforce the market liquidity condition $t_{b}+t_{f}\leq t_{m}$, which guarantees that at least $(1-t_{m})q$ tokens are held by others, creating a pool of market liquidity. Relevant to both purposes, nothing in the bidding mechanism itself creates an incentive for a true report, and, in fact, the incentive is to cheat by hiding some of the toehold to artificially inflate the bid, $A$.
@@
Could a bounty system be effective and code feasible? DAO positions appear as a set of public addresses and token quantities. Commercially available technologies to trace and attribute token ownership exist and have a significant degree of effectiveness at what may be feasible cost. [^49] See Yaffe-Bellany (2023) (discussing Chainalysis). But reliance on outside commercial parties inhibits code feasibility. The DAO may have to implement the bounty system via contracts with outside teams. Doing so effectively while the control party has control of the DAO may not be possible. The task is complicated by the possible need to resolve disputes about the veracity of identifications claimed by the bounty hunters. [^50] Dispute resolution is more difficult when the parties do not have accessible on-chain assets at risk. See note 44 supra. An approach such as requiring good faith deposits from bounty hunters might be required to implement a dispute resolution mechanism that is code feasible. And possible hidden ownership presents a major challenge to the effectiveness of bounty systems. [^51] Bounty hunters do have some possible strategies in the face of hidden ownership. Counterparties to the hidden ownership position used for concealment have an incentive to collect the bounty by disclosing the failure to report, potentially earning additional tokens at the expense of the control party without violating the underlying contract. Bounty hunters may angle for a cut by advertising this opportunity broadly and offering assistance. Although a bounty system may be too difficult to implement, it is potentially a very powerful deterrent. Successful bounty recovery is a disaster for the winning bidder, involving loss not only of potential added token value, the motivation for concealment, but also of the base value, $P_{0}$.
One set of approaches to address control party position reporting issues centers around using know-your-customer registration systems that make token holders identifiable. Requiring registration of all token positions at all times would be very costly both initially and in the face of continual revisions as tokens are traded between parties. Registration might also raise privacy considerations requiring costly zero-knowledge proof or other technologies to make identities private yet verifiable. Nonetheless, use of registration limited to subsets of token holders and to particular points in time might be useful and cost-effective as we describe in what follows.
@@
    Registration. Parties can register through a know-your-customer process to demonstrate that they are not the control party or related to the control party. [^52] Registration can be made consistent with privacy through approaches such as requiring a zero-knowledge proof confirmation that they are not a restricted party linked to the winning bidder. For example, Rosenberg et al. (2022) describe a zero-knowledge proof credentialing system that they call “zk-creds.” In one instantiation, parties embed their passport in a zero-knowledge privacy layer that allows proof of identity. All registered parties are eligible to participate in the token auction. The set of $T_{1}$ token holders who register will be eligible to receive surplus from the auction.
@@ 4.1.6 Auction Efficiency and Market Liquidity
Like second price auction frameworks in other contexts, the English auction that embodies the basic mechanism here has strong efficiency properties. The bidder who can produce the most social surplus wins, with appropriation of that surplus by the winning bidder limited by the level of the second highest bid. As discussed, there is another limitation: The mechanism is subject to the market liquidity constraint. This constraint limits the added token value available to the winning bidder, reducing it by $F=(1-t_{m})(V^{*}-P_{0})q$, the minimum amount that must accrue to free riders. Consider the project $(V^{*},C^{*})$ that creates the highest possible social surplus $\psi^{*}=N^{*}-C^{*}$ and produces total added token value $N^{*}=(V^{*}-P_{0})q$. If $\psi^{*}=N^{*}-C^{*}>0>N^{*}-C^{*}-F$, then this best possible project will not be implemented because it cannot cover its cost and also distribute $F$ to the free riders.
@@
There is another possible source of inefficiency, which we will call “toehold overbidding.” Burkart (1995) shows that in an English auction, it may be optimal for participants with toeholds to bid higher than their valuations. Overbids create a danger of overpaying if the overbidder wins, but they also may push up the winning auction price, benefiting the toehold position if the overbidder loses. Burkart (1995) proves that, in the context of his model, for sufficiently small overbids, the potential toehold benefit outweighs the danger of loss from winning the auction with an overbid. [^54] The potential gain on the toehold is of the order of magnitude of the overbid while the potential loss from the overbid is of the order of magnitude of the overbid squared. Burkart (1995) assumes a private values setting in his main inquiry. Bulow et al. (1999) show that toehold overbidding tends to be a much more serious problem in common values situations.
The mechanism developed here eliminates the incentive to make toehold overbids. Bidders are competing to offer as much added token value to the other token holders as possible. There two ways to offer more added token value: through a higher value claim or a lower surplus claim. However, as long as the value deposit forfeit function satisfies the inequality condition in Lemma 1, bidders are better off increasing their bids by lowering their surplus claims. But doing so has no impact on the corresponding value claims and therefore cannot cause competing bidders to earn more on each share of their toehold positions by overbidding. [^55] The result in Lemma 1, that making excessive value claims is a dominated strategy, remains intact.
Nonetheless, there is a strategy analogous to toehold overbidding that can have similar effects under the mechanism: surplus claim underbidding, lowering the surplus claim below the point that causes the best business plan of the bidder to break even. The goal is to cause the winning bidder to also lower its surplus claim. That smaller surplus claim reduces the proportion of $T_{1}$ token holdings that must be sold to the winning bidder at $P_{0}$ and increases the corresponding proportion on which the $T_{1}$ token holders, including the bidder who engaged in surplus claim underbidding, can earn the full amount of added token value.
@@
Successful surplus claim underbidding is socially desirable because more social surplus is shifted to the $T_{1}$ token holders, which results in a positive investment impact while not affecting the result that the best business plan is put into effect. Unsuccessful surplus claim underbidding, however, has negative social properties. The winning bidder implements an inferior business plan. To the extent of the underbid, additional apparent social surplus flows to the $T_{1}$ token holders based on the value deposit forfeit function, which will result in a positive investment impact. However, this impact is larger than is optimal, because at least part of the surplus is apparent rather than real, being in excess of what would be generated even by the best business plan. [^56] Investment is over-encouraged if we look at this auction event in isolation. It is possible that the transfer in excess of actual social surplus generated by this type of event offsets the failure to shift the full amount of social surplus to the $T_{1}$ token holders in other auctions. In short, the transfer in excess of social surplus may be second-best optimal in some instances. It is not clear how these two impacts balance out across auctions, but it is clear that a net negative effect is possible.
@@ 5.1 Impact on Investment
Another consideration is that sealed-bid first price auctions may avoid the collusion problems that can arise in second price auctions. Collusion coalition parties in a sealed-bid first price auction can defect by bidding above an agreed upon low price chosen to enable the highest valuing bidder to prevail at that price and thus collect the largest possible amount of added token value for the coalition. In a second price auction, the highest valuing bidder will submit their best bid, and other coalition members cannot win the auction by deviating. Deviating only reduces the joint added token value for the coalition. The potential considerations in choosing among auction forms are myriad, and we do not attempt a detailed discussion here. [^64] A good starting point is Krishna (2010). Instead, we simply rest on a claim that given current knowledge, the English auction is a strong candidate with respect to efficiency and raising revenue compared to other standard auctions.656565 We have been assuming a “private values” setting with asymmetric information in which bidder business plans are known only to each bidder and bidder valuations of the plans are independent. See note 39 supra. If, instead, there is a common values element in which bidders have different signals concerning the value of the current operation and bid based on perceived undervaluation at price $P_{0}$, then the efficiency and surplus-distribution of various auction methods are potentially different than under the private values assumption. However, the English auction performs relatively well in a common values setting. Under certain assumptions such as the average crossing property, the English auction remains efficient in the sense that there is an ex-post equilibrium in which the bidder who can add the most value will win. (Krishna, 2010, pp. 134-143). In this setting, the English auction tends to raise more revenue than other approaches, including sealed-bid first price and second price auctions. (Krishna, 2010, pp. 97-100). In the context here, raising more revenue equates to extracting more added token value from the winning bidder, one of our desiderata.
@@ 5.3 Undervaluation and Treasury Raid Scenarios
Potential “treasury raid” scenarios provide an interesting perspective on how the mechanism would operate. DAOs typically have treasuries consisting of cryptocurrency tokens that can be used for further development of the DAO. Treasuries can store retained earnings as well as funds received from investors. Treasuries may be under the direct control of token holders who can direct their use by voting on proposals. [^71] In some cases, there are other parties who have that formal power or a veto on token holder proposals for spending treasury resources. In one common organizational structure, the DAO is embodied as a Swiss Foundation. The Foundation acts a wrapper, creating legal personhood for the DAO, with concomitant limited liability. Under Swiss law, the Foundation must have formal legal control over the DAO treasury. Although the Foundation may in practice typically defer to token holder votes, there is the power to veto vote outcomes if they involve disbursing treasury funds. Direct control by token holders creates the danger of “treasury raids.” The most pernicious example of a treasury raid is one in which a party gains control of the DAO and then transfers the entire treasury to itself by voting positively on a suitable proposal. The party is stealing the other token holders’ stakes, and the term “raid” along with its negative connotations is apt.
@@
First, there may be an expectation that the parties currently guiding the DAO, although good-intentioned, will engage in foolish business plans, gradually wasting the treasury assets. The foolish business plans may be the result of the exhaustion of the original guiding purposes of the DAO. In this case, the social optimum is to end the DAO and distribute the treasury to the token holders, permitting re-investment of the treasury assets in more promising projects. In a situation of capital market adequacy, there will be “vulture investors,” who facilitate this result via a takeover or through activism in the face of reluctant guiding or control parties.727272 Vulture investors already exist in the cryptocurrency space. The saga of Rook DAO, recounted by Gilbert, Aleksandar (2023) and Nelson (2023), is an example. Rook DAO encountered an internal conflict and was trading below the value of its treasury. Co-founders and members of the core team proposed shifting 75% of the tokens to themselves to continue the project, leaving 25% to the rest of the token community. Activist investors facilitated an increase in the community’s share to 60%. Liu (2023) details the claimed role of Arca, one of the activist investors. As described in Arca (2023) and Dorman (2020), Arca’s operations include some that resemble a hedge fund that both makes portfolio investments and engages in activism.
@@
If, instead, the token value was underwater because the guiding parties have a superior business plan that requires secrecy, capital market adequacy means that these parties can obtain funding in confidence to outbid parties intent on liquidation. The other token holders receive at least their pro rata share of the treasury and possibly also some of the social surplus from the envisioned superior business plan. If the guiding parties think they have a superior business plan but cannot convince any funding party of that opinion, then upon liquidation, the guiding parties at least receive their share of the treasury, possibly amounting to enough to implement their business plan through a new project. [^74] The outcome in the Rook DAO events described in note 72 supra appears to represent these kinds of splits between parties intent on implementing a business plan and parties who want to exit through a liquidation that brings the DAO value up to the level of the treasury. The Rook DAO treasury assets were divided between two such groups of parties. The bidding competition central to the mechanism creates a way to divide the treasury assets more cogently, with the key question being whether the parties who want to continue can outbid the liquidation value. If they can, the other parties receive at least the pro rata treasury value of their tokens and likely more. If they cannot, they receive their share of the treasury and can pursue the project with those funds.